Someone please help me
The sale price of a car is $12,590 which is 20% of the original price. What is the original price?


also..
Jim is able to sale a statue for $670 he made a 35% profit. What is the original price?
Someone please help me ill be really appreciative

Answer by stanbon(48568)

The sale price of a car is $12,590 which is 20% of the original price. What is the original price?
-----------------
Let the original price be "x".
EQUATION:
0.20x= 12590
x = 12590/0.20 = $62950.00
------------------------------------

also..
Jim is able to sell a statue for $670; he made a 35% profit. What is the original price?
------
Let the original price be "x".
EQUATION:
x + 0.35x = 670
1.25x = 670
x = $536
=================
Cheers,
Stan H.